OpenStudy (anonymous):

Let f(x)=3.5x^2-5.5x. Find the equation of the secant line on the graph f(x) between x1=6 and x2=13. Write your answer in mx+b format.

OpenStudy (anonymous):

y1 = 3.5(6)^2 - 5.5(6) = 93 y2 = 3.5(13)^2 - 5.5(13) = 520 m = \(\frac{520-93}{13-6}=61\) y - 93 = 61(x - 6) y = 61x-273